Problem 31.14 - Enhanced - with Hints Part B The magnetic field of an electromagnetic wave in a vacuum is B,-(2.4pT) sin((1.00 x 107)z-wt), where z is in m and t is in s You may want to review ( pages 889-892) What is the waves frequency? Express your answer to three signiticant tigures and include the appropriate units. Hints For help with math skills, you may want to review Rearrangement of Equations Involving Multiplication and Division f-Value Units Submit My Answers Give Up Part C What is the waves electric field amplitude? Express your answer to two significant figures and include the appropriate units. Hints EValue Units Submit My Answers Give Up
Answer #1

wave speed c = 3*10^8 m/s


wavenumber k = 2*pi/lambda = 1*10^7


wavelength lambda = 2*pi/(1*10^7) = 6.28*10^-7 m

frequency f = c/lambda = 3*10^8/(6.28*10^-7) = 4.78*10^14 Hz


=================================

part C


Eo = Bo*c


Bo = 2.4*10^-6 T

Eo = 2.4*10^-6*3*10^8


Eo = 720 V/m